# Effective SAM Combination for Open-Vocabulary Semantic Segmentation

### v1

@@ 1 Introduction
The proposed ESC-Net is evaluated on standard open-vocabulary benchmarks, including the ADE20K [39], PASCAL-VOC [9], and PASCAL-Context [25] datasets, achieving state-of-the-art performance compared to previous methods. Additionally, we demonstrate through various ablation studies that our model maintains robust performance across a range of challenging scenarios.
@@ 2 Related Work
SAM’s zero-shot proficiency draws comparisons to the advancements in natural language processing seen in models like ChatGPT, marking it as a milestone toward artificial general intelligence in visual tasks. Prior to SAM, foundation models in computer vision, such as CLIP and ALIGN, focus on aligning multimodal data—specifically text and image encoders—through contrastive learning for cross-modal zero-shot generalization. However, these earlier models primarily excel at tasks reliant on cross-modal alignment, such as image captioning and generation, rather than segmentation. SAM advances beyond these models by specifically targeting image segmentation as a foundational task, making it a versatile tool for addressing complex segmentation challenges.

### v2

@@ 1 Introduction
The proposed ESC-Net is evaluated on standard open-vocabulary benchmarks, including the ADE20K [39], PASCAL-VOC [9], and PASCAL-Context [25] datasets, achieving state-of-the-art performance compared to previous methods. Additionally, we demonstrate through various ablation studies that our model efficiently maintains strong performance across a range of challenging scenarios.
@@ 2 Related Work
Prior to SAM, foundation models in computer vision, such as CLIP and ALIGN, focus on aligning multimodal data—specifically text and image encoders—through contrastive learning for cross-modal zero-shot generalization. However, these earlier models primarily excel at tasks reliant on cross-modal alignment, such as image captioning and generation, rather than segmentation. SAM advances beyond these models by specifically targeting image segmentation as a foundational task, making it a versatile tool for addressing complex segmentation challenges.
@@ 4.3 Implementation Details
In traditional open-vocabulary segmentation methods [10, 6, 35], the image-text relationship modeling of pre-trained CLIP is typically preserved by freezing its encoders. However, recent studies have demonstrated that fine-tuning CLIP’s image and text encoders can achieve higher performance in open-vocabulary segmentation tasks. This suggests that CLIP’s linguistic discrimination capability can be retained even with training datasets containing a limited number of classes. Following this insight, the proposed ESC-Net adopts the approach of CAT-Seg [5] to fine-tune the attention layers of CLIP’s image and text encoders.
@@ 5 Conclusion
Acknowledgement. This work was supported by the Korea Institute of Science and Technology (KIST) Institutional Program (Project No.2E33001-24-086), National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(No. RS-2024-00423362), and Yonsei Signature Research Cluster Program of 2024 (2024-22-0161).